OpenStudy (anonymous):

use the double angle formula to find the exact value of sin(2arccos(4/5)?

OpenStudy (anonymous):

i understand that u let arccos(4/5) = theta but how do i find the arccos(4/5)

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }=\cos \theta \]  \[\sin \theta=\frac{ 3K }{ 5K }=\frac{ 3 }{ 5 }\] \[\sin 2\theta=2 \sin \theta \cos \theta \] \[=2*\frac{ 3 }{5 }*\frac{ 4 }{5 }=\frac{ 24 }{25 }\]